OpenStudy (anonymous):

A Can Anyone Help? 5 foot-long cylindrical pipe has an inner diameter of 6 feet and an outer diameter of 8 feet. If the total surface area is kn(n=Pi) ,what is the value of k ?

OpenStudy (anonymous):

Area = Pi*6*5 + Pi*8*5 = Pi(6*5+8*5)=70 Pi k= 70

OpenStudy (anonymous):

@Identified